OpenStudy (anonymous):

Rewrite the radical as a rational exponent. the cube root of 2 to the seventh power

OpenStudy (anonymous):

\[\huge x ^{m/n} = \sqrt[n]{x^m} = (\sqrt[n]{x})^m\]

OpenStudy (anonymous):

That's it iambatman just said everything I was about to say. Use that formula and you will be able to get an answer.

OpenStudy (anonymous):

so 2 7/3 ?

OpenStudy (kainui):

So how do we make roots as exponents? First off if you have something like \[\LARGE (n^3)^2=(n^3)(n^3)=(n*n*n)*(n*n*n)=n^6\] That tells us, ahh! It looks like the rule is to multiply exponents if they're something exponentiated raised to an exponent. So now we look at \[\LARGE \sqrt{x}^2=x^1\] So it looks like the x^2 is cancelled out by the square root to equal 1. This is just like the last case, so it must mean that \[\Large \sqrt{x}=x^{1/2}\] And this makes sense that it will be true for other roots too. Cube root should cancel out with the cube to be 1, so that should also be 1/3 power for cube root.

OpenStudy (anonymous):

\[2^{7/3}\]

OpenStudy (anonymous):

and to do this reversed? Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power

OpenStudy (anonymous):

how would I do that?

OpenStudy (anonymous):

@WordGEEK @iambatman @Kainui

OpenStudy (anonymous):

(53/4) 2/3

OpenStudy (anonymous):

What are you trying to ask now?

OpenStudy (anonymous):

what would (53/4)2/3 be as a radical ?

OpenStudy (aum):

\(\Large (5^{ \frac 34})^ \frac 23 \) ?

OpenStudy (anonymous):

yes^ :)

OpenStudy (aum):

Since \(\Large (x^m)^n = x^{mn}\) Multiply the exponents first. What do you get?

OpenStudy (anonymous):

6?

OpenStudy (aum):

\(\Large \frac 34 * \frac 23\) = ?

OpenStudy (anonymous):

1/2 ?

OpenStudy (anonymous):

Yup good it equals 1/2 so now just apply the formula that iambatman posted.

OpenStudy (aum):

Yes. \(\Large x^{\frac 12} = \sqrt{x}\). Therefore, \(\Large 5^{\frac 12}\) = ?

OpenStudy (anonymous):

2\[\sqrt[2]{5}\]

OpenStudy (anonymous):

and the one on the outside ?

OpenStudy (anonymous):

I don't think that's right

OpenStudy (aum):

Yes. But 2 is a special case where \(\Large \sqrt [2]{5} = \sqrt{5}\) Both mean the same. Square root 5 or just radical 5. \(\Large (5^{ \frac 34})^ \frac 23 = 5^{\frac 34 * \frac 23} = 5^\frac 12 = \sqrt {5}\)

OpenStudy (anonymous):

oh okay! now it makes sense! that's one of the answer options. thank you! :)

OpenStudy (aum):

You are welcome.